OpenStudy (mathhelp346):

3x-(2+6x) less than or equal to 28 i keep getting 10/3? did i do something wrong?

OpenStudy (anonymous):

-3x<=30 x<=30/-3 x<=-10 is what I got

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

to you?

OpenStudy (mathhelp346):

um no

OpenStudy (anonymous):

\[3x-2-6x\] \[-3x-2\]

OpenStudy (anonymous):

can you see that now?

OpenStudy (anonymous):

that's the leftside of your equation. add 2 to both sides. and then divide both sides by -3 then you will get -10 as your answer

OpenStudy (mathhelp346):

ut wouldn't u add -6x to -6x and 3x?

OpenStudy (mathhelp346):

i mean add 6x

OpenStudy (anonymous):

when you see -(2+6x) this means you distribute the negative to both terms so -2-6x is what it looks like with the brackets removed so you have +3x and -6x when you add those 2 terms you get -3x

OpenStudy (mathhelp346):

ohhh ok

OpenStudy (anonymous):

3x-(2+6x)<or equal to 28. => 3x-2-6x<or equal to 28. => -3x-2< or equal to 28. => 3x+2> or equal to 28. (Note the change in sign). 3x>or equal to 26 Therefore x is > or equal to 26/3. This is what I got. -__- What's the correct answer?

OpenStudy (anonymous):

i wouldn't do that.

OpenStudy (anonymous):

Why not? o.o?

OpenStudy (anonymous):

Is it wrong? :/

OpenStudy (anonymous):

http://www.webmath.com/_answer.php

OpenStudy (anonymous):

yes, it was wrong but there's how a calculator does it

OpenStudy (anonymous):

The page isn't opening. So, what is the correct answer?

OpenStudy (anonymous):

well technically i made a mistake! so actually when you divide by the negative you should get as a final answer x>=-10

OpenStudy (anonymous):

and that is correct

OpenStudy (anonymous):

3x-(2+6x)<=28 =3x-2-6x<=28 =-3x-2<=28 +3x +3x =-2<=28+3x =-28 -28 =-30<=3x =-10<=x

OpenStudy (anonymous):

Yup. Got it. Lol. That was such a horrible mistake on my part. I forgot to put the -ve sign on 28. :/ Thanks. (:

OpenStudy (anonymous):

be careful with the <> symbol. I need to correct mine. It should be -10>=x